OpenStudy (anonymous):

Use the five steps for problem solving to answer the following question. Please show all of your work. The area of a triangular section of a store cannot be more than 910 sq. ft. If the height of the triangular section is 52 feet, what should the base be?

OpenStudy (anonymous):

the area of a triangle is \[a =(base*height)/2\] if you already have height, just do the following: \[base = (2*area)/height\]

OpenStudy (anonymous):

so \[A \le 910\] A = 1/2*b*h so \[1/2*b*h \le 910\] since h = 52ft. 1/2*b*52 <= 910. 26b<= 910 b<=35. :)

OpenStudy (anonymous):

ok im in agreement whats the arrows for is that representing the equal sign